The lifetime T in tens of hours of a battery has a cumulative distribution function

F(t)={0,t<1

{k(t2+2t-3),1<t<or equal to 1.5

{1,t>5

Find k and hence find p(T>1.2) and find the median of T


1
Expert's answer
2021-07-13T10:40:55-0400
"F(t) = \begin{cases}
 0, & t<1 \\
 k(t^2+2t-3), &1\leq t\leq 1.5 \\
 1, & t>1.5
\end{cases}"

1. Differentiate with respect to "t"


"f(t) = \begin{cases}
 
 k(2t+2), &1\leq t\leq 1.5 \\
 0, & otherwise
\end{cases}"

"\displaystyle\int_{-\infin}^{\infin}f(t)dt=\displaystyle\int_{1}^{1.5}k(2t+2)dt"

"=k[t^2+2t]\begin{matrix}
 1.5 \\
 1
\end{matrix}=k(2.25+3-1-2)=2.25k=1"

"k=\dfrac{4}{9}"

2.


"P(T>1.2)=1-P(T\leq1.2)=1-F(1.2)"

"=1-\dfrac{4}{9}(1.2^2+2(1.2)-3)=\dfrac{1.88}{3}\approx0.6267"

3.


"\dfrac{4}{9}(m^2+2m-3)=0.5"

"m^2+2m-3=1.125"

"m^2+2m-4.125=0"

"m=-1\pm\sqrt{5.125}"

SInce median ">0," we take


"median=-1+\sqrt{5.125}\approx1.264"